OpenStudy (anonymous):

Simplify completely:

OpenStudy (anonymous):

\[\frac{ 3x ^{4}+5x ^{2} +2x}{ x }\]

OpenStudy (mertsj):

Break it up into 3 fractions like this: \[\frac{3x^4}{x}+\frac{5x^2}{x}+\frac{2x}{x}\]

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

cancel the common factors and simplify:)

OpenStudy (anonymous):

idk how

OpenStudy (ajprincess):

\[\frac{3x^4}{x}=\frac{3*x*x*x*x}{x}\] \[=\frac{3*\cancel x*x*x*x}{\cancel x}\] \[=3*x*x*x\] \[=3x^3\] Similarly simplify the rest of the fractions.:)

OpenStudy (anonymous):

So the answer would be: \[3x^{3} + 5x + 2\]

OpenStudy (ajprincess):

yup right:)